OpenStudy (anonymous):

find a formula for the derivatives of the function using different quotients m(x)= 1/(x+1)

OpenStudy (anonymous):

y=(x+1)^-1 (actually you can write x=1 to the power -1) dy/dx= -1*(x+1)^-1-1=-1*(x+1)^-2=-1/(x+1)^2